OpenStudy (luigi0210):

Quick derivative question:

OpenStudy (luigi0210):

What is \(\LARGE \frac{d}{dx}~(arctan~4x)\)?

OpenStudy (johnweldon1993):

If my memory is correct it would be \[\large \frac{4}{1 + 16x^2}\] Basically by using the chain rule

OpenStudy (anonymous):

Correct.

OpenStudy (johnweldon1993):

\[\large \frac{d}{dx} tan^{-1}(x) = \frac{1}{1 + u^2}\] \[\large \frac{d}{dx} 4x = 4\] inside times outside = \(\Large \frac{4}{1 + 16x^2}\)

OpenStudy (luigi0210):

Alright, thank you guys :)

sammixboo (sammixboo):

dx/d = something *shrugs*